Friges in an interference pattern is observed to be 2mm apart. if the silt separation is 1mm and the wavelenght of light is 400mm. What is the distance of the screen where the observation was made?

1
Expert's answer
2022-03-02T14:38:38-0500

Distance can be found as follows (see https://www.khanacademy.org/science/physics/light-waves/interference-of-light-waves/v/youngs-double-slit-problem-solving):


"D = \dfrac{yd}{\lambda}"

where "\lambda = 400nm = 400\times 10^{-9}m", "y = 1mm = 10^{-3}m" and "d = 2mm = 2\times 10^{-3}m". Thus, find:


"D= \dfrac{2\times 10^{-3}m\cdot 10^{-3}m}{400\times 10^{-9}m} =5m"

Answer. 5m.